You have $1,000 and you have 2 investment options: A) You can invest all of your money in an account that pays 5%/year or B) You can invest half of your money in an account that pays 8%, but the other half will sit in your checking account earning nothing. If you plan to leave your money in these accounts for 7 years, which one is the better investment and how much money do you have after 7 years?
Answer #1

Amount = 1000$

Option 1 : when you invest all money in a account that pays 5% annually

Amount = 1000$

Time (N) = 7 Years

Interest rate = 5% annually

Future worth = Present value (1 + Rate of interest) ^n

= 1000 ( 1 + 5%)^7

= 1000(1 + 0.05)^7

= 1000 (1.05)^7

= 1000 * 1.41 = 1,410 $ in 7th year

Option 2 : When half of money is invested at 8 % and rest kept in checking or current account that pays no interest

Amount = 500$

Time = 7 Years

Interest rate = 8% annually

Future worth = Present value (1 + Rate of interest) ^n

=500 (1 + 8%)^7

= 500(1+0.8)^7

= 500(1.08)^7

= 500 * 1.71 = 855 $

So, total amount at the end of 7th year = 855 + 500(checking acc)

= 1355$

So, here 1st option is better than 2nd option for investment purpose